Sequences Applications: Stadium Construction
A stadium has 2040 seats. The first row has 10 seats and each successive row has 4 additional seats. How may rows are there in the corner section?

Answer by venugopalramana(3286)

Sequences Applications: Stadium Construction
A stadium has 2040 seats. The first row has 10 seats and each successive row has 4 additional seats. How may rows are there in the corner section?
TOTAL SEATS ARE
2040=10+14+18+22+................
THIS IS AN A.P.
SN=2040
N=?
A=10
D=4
SN = (N/2)[2A+(N-1)D]
2040= (N/2)[2*10+(N-1)4]
2*2040=20N+4N^2-4N
4N^2+16N-4080=0
N^2+4N-1020=0
N^2+34N-30N-1020=0
N(N+34)-30(N+34)=0
(N-30)(N+34)=0
HENCE
N=30
THER ARE 30 ROWS.